Using Mohr's circle determine for the below differential element:

a) the principal stresses and the plan on which they act. Show the stresses on a properly oriented differential element. Label all stresses.

b) the maximum shear stress and the plan on which they act. Show the stresses on a properly oriented differential element. Label all stresses.

c) the stresses on a differential element 40 degrees clockwise from the original element. Show the stresses on a properly oriented differential element. Label all stresses.

Sigma X = -20 (ksi) Tau XY = 15 (ksi) Sigma Y = 5 (ksi)

The maximum principal stress on Mohr's circle is _____ (ksi)

The minimum principal stress on Mohr's circle is ______ (ksi)

The angle at which the principal stresses occurs is ______ (degrees- counter clockwise)

The maximum shear stress on Mohr's circle is ______ (ksi)

The angle at which the maximum shear occurs is ______ (degrees - clockwise)

бу ту yx тху Ox ох Txy тух ТЕ Oycan someone help me with this mohr's circle problem?
Answer #1

6 eng eng on .br gy n=-20ksi Uny = isksi Gy sksi C= Caug= Caug= bn toy -20+5 2 -7.5ksi Radius = lsnca) + (May)? V (+20,25) +shear stress Principal stress tan 20 = 4ny Ontby 20, -50.189 285 max = 50-18+90 = 140.18° tan 28, 2x15 25 Osmax = 70.09° 0 04